OpenStudy (anonymous):

Find the solution of 3 square root of x+5=-9 these are the anser choices: x = 4; not extraneous x = 4; extraneous x = 0; extraneous x = 0; not extraneous

OpenStudy (dls):

can you write the q in a better way perhaps?

OpenStudy (anonymous):

\[\sqrt[3]{2}\]

OpenStudy (anonymous):

my bad hold up

OpenStudy (anonymous):

\[\sqrt[3]{x2}\]

OpenStudy (anonymous):

omgsorry i messed up

OpenStudy (anonymous):

hold up i got it

OpenStudy (dls):

\[\LARGE 3 \sqrt {x+5} = -9\] ?

OpenStudy (anonymous):

3\[3\sqrt{x+5=-9}\]

OpenStudy (anonymous):

yeah thats it! sorry im new to this

OpenStudy (dls):

the question is wrong..sqrt always gives a +ve number..how can RHS be negative still..square both the sides \[\LARGE 9 (x+5) = 81 = x+5 = 9 => x=4\] but that doesn't satify the equation..it does if RHS = + 9

OpenStudy (anonymous):

thats what the question said

OpenStudy (dls):

x=4,extraneous is the answer

OpenStudy (anonymous):

ok thank veryy much